A gold processor has three sources of gold ore, Alaska California, and Utah. Table below shows gold ore details: (20 points) Alaska California Utah Cost per ton [$] $180 $120 $100 Gold content [g/ton] 4 3.5 3.2 Processing time [min/ton] 7 5 5 The gold processor wants to maximize its gold production. The processing plant works from 8:00 to 20:00. The gold processor can spend $14,000 per day in raw materials. California environmental regulations limit daily ore material extraction to 45 tons. Finally, a contract states that the amount of ore used from Alaska has to be at least 50 tons. Formulate a linear programming problem to maximize the gold production.
Answer #1

Linear programming model is following:

Decision variables:

Let

X1 = tons of ore used from gold ore in Alaska daily

X2 = tons of ore used from gold ore in California daily

X3 = tons of ore used from gold ore in Utah daily

Objective function:

Max 4X1+3.5X2+3.2X3 (total daily gold production)

s.t.

Constraints:

X2 <= 45 (ore extracted from california is limited to 45 tons)

X1 >= 50 (ore extracted from Alaska has to be at least 50 tons)

180X1+120X2+100X3 <= 14000 (daily budget for raw materials)

7X1+5X2+5X3 <= 720 (12*60 = 720 minutes from 8:00 to 20:00)

X1, X2, X3 >= 0